SpaceRef

SpaceRef


NOAA GOES SATOPS Morning Report - August 7, 2003† Day 219†

Status Report From: National Oceanic and Atmospheric Administration
Posted: Thursday, August 7, 2003

DMSP†

OPS49 rev 18855 R+10 from POGO-C at 2032Z on August 6:† SCCR# 49-37 completed for a SCP 1 clock drift correction.† On rev 18856 at 2212Z, the clock drift correction was applied to SCP 2.†

All DMSP operations were nominal over the past 24 hours.†

POES†

NOAA-17 rev 05804 / F at 1854Z on August 6:† A 13M 'B' antenna failure at AOS necessitated failover to the 'A' antenna.† The failover was accomplished with minimal data loss.† All of GAC2A, 77% of LAC1A, and 91% of LAC4B were received and recorded.† All of the post-pass transfers were delayed for more than an hour beyond their scheduled transfer times due to other higher priority commitments for the antenna and CDA personnel.† Recorder management prevented any replays from the spacecraft.†

GOES†

At 1348Z on August 6, a GOES-9 frame break was observed and attributed to high BER (unknown).†

At 1434Z and 1518Z on August 6, GOES-10 frame breaks were observed and attributed to high BER (unknown).†

GOES-8 continues its westward drift at 1.06 deg/day.† The drift table with the estimated current position can be found at† http://www.oso.noaa.gov/goes/GOES-8_Drift_Table.txt.† GOES-8 will be stopped at 165E on or about August 24.†

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter